# Fine Mixed Subdivisions of a Dilated Triangle

### v1

@@ Abstract.
An upward equilateral triangle of side $n$ can be partitioned into $n$ unit upward equilateral triangles and $\frac{n(n-1)}{2}$ unit rhombi with $60^{\circ}$ and $120^{\circ}$ angles. In this paper, we focus on understanding such partitions with a fixed arrangement of unit triangles. We formulate a criterion for such a partition being unique, identify a set of operations that connects all such partitions, and determine which parts are common to all such partitions. Additionally, we discuss an operation that connects all possible partitions with different arrangements of triangles.
@@ 1. Introduction
Consider an equilateral triangle of side $n$. Suppose that we cut out $n$ upward unit equilateral triangles, and attempt to tile the rest with unit 60-120 rhombi (formed by gluing two unit equilateral triangles together). This results in a fine mixed subdivision of the triangle, sometimes also called a rhombus tiling of a holey triangle ([1]).
@@
In [1], Ardila and Billey described the condition of the positions of the holes for which the tiling is possible via a “spread-out condition” (see Theorem 2.3). In this paper, we focus on extending our understanding of the tilings with any given set of hole positions beyond their existence. Sections 3-6 explore the structure of the set of tilings for a fixed arrangement of triangular holes, and Section 7 discusses an operation that moves the holes and connects all possible tilings.
@@ 2. Preliminaries
A fine mixed subdivision of an upward equilateral triangle of side length $n$ is a partition into $n$ upward unit equilateral triangles (holes) and $\binom{n}{2}$ unit 60-120 rhombi.
(From now on, we will simply use “rhombus” to refer to a unit 60-120 rhombus. We will also drop the “equilateral” adjective for triangles as all triangles in this paper will be equilateral.)
Suppose we’re given a set of triangular holes and what to find out whether there is a fine mixed subdivision with this set of holes. Note that every upward triangle $T$ of side $k$ consists of $k$ more upward unit triangles than downward unit triangles. Also, note that a rhombus covers an equal number of upward and downward triangles in $T$ if it does not intersect the boundary of $T$, and covers one upward triangle if it intersects the boundary of $T$. Hence, there are at most $k$ triangular holes that lie in $T$. Turns out, if this condition is satisfied for every $T$, then the desired fine mixed subdivision exists.
@@
We will say that an arrangement of $n$ upward unit triangular holes in an upward triangle of side $n$ is spread-out if every upward grid triangle of size $k$ contains at most $k$ holes. We will also refer to the side length of such a triangle as its size.
@@
An arrangement of holes is a part of some fine mixed subdivision if and only if it’s spread-out.
@@ 3. Uniqueness of Tiling
More precisely, consider an upward triangle $\Delta$ of side length $n$ divided into $n^{2}$ unit triangles, and suppose that we cut out several disjoint upward triangular holes whose sides sum to $n$.
@@
Similarly to the case of unit triangular holes, one may prove that if $\Delta\setminus\cup_{s\in S}s$ is tileable with rhombi, then $S$ is spread-out. If we replace all triangles in $S$ with upward unit triangles in their bottom rows, we get a spread-out arrangement of unit triangles, so $S$ being spread-out is sufficient for $\Delta\setminus\cup_{s\in S}s$ to be tileable with rhombi.
For upward grid triangles $s_{1}$, $s_{2}$, denote by $s_{1}\vee s_{2}$ the minimal upward triangle that contains both $s_{1}$ and $s_{2}$. Note that if saturated triangles $s_{1},s_{2}\in S$ intersect at exactly one point, then $s_{1}\vee s_{2}$ is saturated, and there is only one way to tile $(s_{1}\vee s_{2})\setminus(s_{1}\cup s_{2})$. Also, no other triangles in $S$ can intersect $s_{1}\vee s_{2}$ in their interiors. Therefore, there is an obvious bijection between tilings of $\Delta$ with holes $S$ and tilings of $\Delta$ with holes $S\setminus\{s_{1},s_{2}\}\cup(s_{1}\vee s_{2})$.
@@
Suppose that we start with some arrangement of holes and begin performing changes of the following type: two triangles $s_{1}$, $s_{2}$ touching at a point can be replaced with $s_{1}\vee s_{2}$. Eventually, we will get an arrangement of triangular holes that are pairwise not touching. If this arrangement consists of a single hole coinciding with the big triangle, then the original arrangement of holes has a unique tiling. We claim that the converse is also true.
@@
This is sequence A295928 on OEIS, which counts the number of triangular matrices $\{T_{i,j}\mid 1\leq i,j\leq n,i+j\leq n+1\}$ with exactly $n$ $1$-entries (and the rest being 00’s) such that one can repeatedly replace triples of the form $(T_{i,j},T_{i+1,j},T_{i,j+1})$ with exactly two $1$’s with all $1$’s until the entire matrix only contain $1$’s. It is not difficult to see that this description is the same as our condition.
@@
Based on the uniqueness condition, it suffices to show that if two triangles $s_{1}$ and $s_{2}$ touch and each has a regular subdivision, then $s_{1}\vee s_{2}$ also has a regular subdivision. We use a different condition of regularity that applies to fine mixed subdivisions of dilated triangles:
Lemma 3.6 ([3], Theorem 1, paraphrased).
From Lemma 3.6 we can assume that $s_{1}$ and $s_{2}$ each correspond to some tropical line arrangement $T_{1}$ and $T_{2}$. Assume without loss of generality that $T_{2}$’s top vertex is touching $T_{1}$’s bottom edge, as shown in Figure 4. In this case, we can shrink $T_{2}$ and place all of its vertices between the two downward rays of $T_{1}$ that correspond to the two sides of the point $s_{1}\cap s_{2}$, and sufficiently far below so that the leftward and rightward arms of $T_{2}$ are all below $T_{1}$’s. It is not difficult to see that this gives a realization of the unique tiling of $s_{1}\vee s_{2}$. ∎
@@ 4. A Graph Representation of Fine Mixed Subdivision
Consider some fine mixed subdivison. Note that it divides the plane into regions: triangular holes, rhombi, and the outside region. Consider the graph $G$ whose vertices are these regions, and where there is an edge $a\rightarrow b$ if and only if $a$ contains an upward triangle adjacent to a downward triangle contained by $b$. We say that $G$ is the graph corresponding to the original fine mixed subdivison.
@@
Note that that all vertices corresponding to rhombi have indegree and outdegree equal to $2$; each vertex corresponding to a triangular hole of side $k$ has outdegree $3k$ and indegree 00; and the vertex corresponding to the outside region has outdegree 00 and indegree $3n$.
@@
Consider a spread-out arrangement where no two trianglular holes touch, and any rhombus $a$ of some tiling. We claim that there is at least one rhombus $b$ such that the graph contains the edge $a\leftarrow b$. Indeed, each of the two edges going into $a$ is coming from either a triangle or a rhombus; if both come from a triangle then the two triangles are touching. Thus, we can construct an infinite sequence or edges $a_{1}\leftarrow a_{2}\leftarrow a_{3}\leftarrow\dots$, so there must be a cycle in the graph, and we can perform a cycle flip, giving us a second tiling. ∎
@@ 5. GD Flips
Using the the idea from the proof of Lemma 4.1, we see that any two fine mixed subdivisions with the same arrangement of triangular holes differ by several cycle flips. Additionally, if we place $3$ unit triangles in the corners and one triangle of side $n-3$ in the middle we get an arrangement of triangular holes that admits only two fine mixed subdivisions, where only one cycle flip is possible. (See Figure 7.)
@@
For $n\geq 3$, consider the region obtained by removing three corner unit triangles from a triangle of side $n$, as well as removing a triangle of side $n-3$ in the middle. There are two configurations of rhombi that tile this region. Let the Clockwise GD (CW-GD) of size $n-3$ be the configuration where vertical rhombi (the ones without horizontal edges) go along the left side of the triangle, and let Counter-clockwise GD (CCW-GD) be the other one. [^1] This name is inspired by the (pre-2020) Google Drive logo. A GD flip consists of replacing a CW-GD with a CCW-GD. In particular, a GD flip is a cycle flip.
@@
First, we claim that there are no edges going from a rhombus strictly inside $C$ to a rhombus of $C$. Suppose there is such an edge $a\rightarrow b$. Without loss of generality, we may assume that this the boundary between $a$ and $b$ is horizontal, that is, $a$ is directly below $b$. Then we can construct a sequence of rhombi starting with $b,$ $a$ where each next rhombus is directly below the previous one. Eventually, this sequence must hit $C$ again. Thus, we’ve constructed a path surrounded by $C$ connecting two vertices of $C$. This path creates a clockwise cycle smaller than $C$, which is a contradiction.
@@
Suppose that we begin with some tiling and start performing GD flips while possible. Since the number of vertical rhombi does not change after a GD flip, and the sum of $x$-coordinates of their centers decreases with every GD flip, we must eventually stop. Note that there is at most one tiling with no clockwise cycles: if there were two, they would differ by several cycle flips, so one of them would have a clockwise cycle. This shows that starting from any tiling, we can get to the unique tiling with no clockwise cycles by performing GD flips, and then back to any other tiling by performing their inverses. ∎
@@ 7. Trapezoid Flips
For any triangle hole $s$, let $h_{1}(s),$ $h_{2}(s)$, and $h_{3}(s)$ denote the distances from sides $s$ to the parallel sides of the big triangle. Note that for any triangle $s$ with $h_{i}(s)>0$, there is a trapezoid flip that decreases $h_{i}(s)$ by $1$. Then the distance from a fine mixed subdivision with triangle set $S$ to the unique fine mixed subdivision with all triangles at the bottom is at most $\sum_{s\in S}h_{i}(s).$ Then the distance between two fine mixed subdivisions with triangle sets $S_{1}$ and $S_{2}$ is at most $\sum_{s\in S_{1}}h_{i}(s)+\sum_{s\in S_{2}}h_{i}(s).$ Since for any $s$ we have $h_{1}(s)+h_{2}(s)+h_{3}(s)=n-1,$ the sum of this expression over $i=1,2,3$ is equal to $2n(n-1)$, so for some choice of $i$ its at most $\frac{2n(n-1)}{3}$, as desired.
@@ 8. Further Questions
The rhombus tilings of a hexagon with sides $a,b,c,a,b,c$ are in bijection with $3$-dimensional Young tableaux that fit in a $a\times b\times c$ box. A GD flip of size 00 in this case is also known as a hexagon flip, and corresponds to removing a cube.
Figure 13. A GD flip of size 00 (note that this is rotated by $60^{\circ}$ counterclockwise relative to conventions in the rest of the paper for ease of visualization).

### v2

@@ Abstract.
An equilateral triangle of side $n$ can be tiled with $n$ unit equilateral triangles and $\frac{n(n-1)}{2}$ unit rhombi with $60^{\circ}$ and $120^{\circ}$ angles. In this paper, we focus on understanding such tilings with a fixed arrangement of unit triangles. We formulate a criterion for such a tiling being unique, identify a minimal set of operations that connects all such tilings, and determine which parts are common to all such tilings. Additionally, we establish a nearly tight bound on the number of “trapezoid flips” needed to connect any two tilings with different arrangements of triangles.
@@ 1. Introduction
Consider an equilateral triangle of side $n$. Suppose that we cut out $n$ upward unit equilateral triangles, and attempt to tile the rest with unit $60^{\circ}$$120^{\circ}$ rhombi formed by gluing two unit equilateral triangles together. This results in a fine mixed subdivision of the triangle, sometimes also called a rhombus tiling of a holey triangle ([1]).
@@
In [1], Ardila and Billey described the condition of the positions of the holes for which the tiling is possible via a “spread-out condition” (see Theorem 2.3). In this paper, we focus on extending our understanding of the tilings with any given set of hole positions beyond their existence. Sections 3 to 6 explore the structure of the set of tilings for a fixed arrangement of triangular holes, and Section 7 discusses an operation that moves the holes and connects all possible tilings.
@@ 2. Preliminaries
A fine mixed subdivision of an upward equilateral triangle of side length $n$ is a partition into $n$ upward unit equilateral triangles (also referred to as holes later) and $\binom{n}{2}$ unit-side-length rhombi formed by combining two unit equilateral triangles edge-to-edge.
(From now on, we will simply use “rhombus” to refer to such a rhombus. We will also drop the “equilateral” adjective for triangles as all triangles in this paper will be equilateral.)
In the rest of the paper, we will sometimes also call such a fine mixed subdivision a rhombus tiling of a holey triangle, where the “holey triangle” is the region formed by removing the $n$ unit triangles from the large triangle. Such a rhombus tiling is naturally defined for any region comprised of several unit triangles on a triangular grid.
Suppose we’re given a holey triangle and want to determine whether it has a rhombus tiling. Note that every upward triangle $T$ of side $k$ consists of $k$ more upward unit triangles than downward unit triangles. Also, note that a rhombus covers an equal number of upward and downward triangles in $T$ if it does not intersect the boundary of $T$, and covers one upward triangle if it intersects the boundary of $T$. Hence, there are at most $k$ triangular holes that lie in $T$. Ardila and Billey showed that if this condition is satisfied for every $T$, then the desired tiling exists.
@@
We will say that an arrangement of $n$ upward unit triangular holes in an upward triangle of side length $n$ is spread-out if every upward grid-aligned triangle of size $k$ contains at most $k$ holes. We will also refer to the side length of such a triangle as its size.
@@
An arrangement of unit triangles is part of some fine mixed subdivision if and only if it’s spread-out.
@@ 3. Uniqueness of Tiling
More precisely, consider an upward triangle $\Delta$ of side length $n$ divided into $n^{2}$ unit triangles, and suppose that we cut out several disjoint upward triangular holes whose sizes sum to $n$.
@@
Similarly to the case of unit triangular holes, one may prove that if $\Delta\setminus\bigcup_{s\in S}s$ is tile-able with rhombi, then $S$ is spread-out. If we replace all triangles in $S$ with upward unit triangles in their bottom rows, we get a spread-out arrangement of unit triangles, so $S$ being spread-out is sufficient for $\Delta\setminus\bigcup_{s\in S}s$ to be tile-able with rhombi, as a corollary of Theorem 2.3.
For upward triangles $s_{1}$, $s_{2}$, denote by $s_{1}\vee s_{2}$ the smallest upward triangle that contains both $s_{1}$ and $s_{2}$. Note that if saturated triangles $s_{1},s_{2}\in S$ intersect at exactly one point, then $s_{1}\vee s_{2}$ is saturated, and there is only one way to tile $(s_{1}\vee s_{2})\setminus(s_{1}\cup s_{2})$. Also, no other triangles in $S$ can intersect $s_{1}\vee s_{2}$ in their interiors. Therefore, there is an obvious bijection between tilings of $\Delta$ with holes $S$ and tilings of $\Delta$ with holes $S\setminus\{s_{1},s_{2}\}\cup(s_{1}\vee s_{2})$.
@@
Suppose that we start with spread-out arrangement of holes and begin performing changes of the following type: two triangles $s_{1}$, $s_{2}$ touching at a point can be replaced with $s_{1}\vee s_{2}$. Eventually, we will get an arrangement of triangular holes that are pairwise not touching. If this arrangement consists of a single hole coinciding with the big triangle, then the original arrangement of holes has a unique tiling. We claim that the converse is also true.
@@
This is sequence A295928 on OEIS, which counts the number of triangular matrices $\{T_{i,j}\mid 1\leq i,j\leq n,i+j\leq n+1\}$ with exactly $n$ 11-entries (and the rest being 0’s) such that one can repeatedly replace triples of the form $(T_{i,j},T_{i+1,j},T_{i,j+1})$ with exactly two $1$’s with all $1$’s until the entire matrix only contain $1$’s. It is not difficult to see that this description is the same as our condition.
@@
In the case of fine mixed subdivisions of dilated triangles, there is an equivalent condition of regularity that is more useful for our purpose.
Lemma 3.5 ([3], Theorem 1, paraphrased).
@@
Based on the uniqueness condition, it suffices to show that if two triangles $s_{1}$ and $s_{2}$ touch and each has a regular subdivision, then $s_{1}\vee s_{2}$ also has a regular subdivision.
From Lemma 3.5 we can assume that $s_{1}$ and $s_{2}$ each correspond to some tropical line arrangement $T_{1}$ and $T_{2}$. Assume without loss of generality that $s_{2}$’s top vertex is touching $s_{1}$’s bottom edge, as shown in Figure 4. In this case, we can shrink $T_{2}$ and place all of its vertices between the two downward rays of $T_{1}$ that correspond to the two sides of the point $s_{1}\cap s_{2}$, and sufficiently far below so that the leftward and rightward arms of $T_{2}$ are all below $T_{1}$’s. It is not difficult to see that this gives a realization of the unique tiling of $s_{1}\vee s_{2}$. ∎
@@ 4. A Graph Representation of Fine Mixed Subdivision
Consider some fine mixed subdivison. Note that it divides the plane into regions: triangular holes, rhombi, and the exterior region. Consider the graph $G$ whose vertices are these regions, and where there is an edge $a\rightarrow b$ if and only if $a$ contains an upward triangle adjacent to a downward triangle contained by $b$, or $a$ is a region adjacent to the boundary and $b$ is the exterior region. We say that $G$ is the graph corresponding to the original fine mixed subdivision.
@@
Note that that all vertices corresponding to rhombi have indegree and outdegree equal to $2$; each vertex corresponding to a triangular hole of side $k$ has outdegree $3k$ and indegree 0; and the vertex corresponding to the outside region has outdegree 0 and indegree $3n$.
@@
Consider a spread-out arrangement where no two trianglular holes touch, and any rhombus $a$ of some tiling of this holey triangle. We claim that there is at least one rhombus $b$ such that the graph contains the edge $a\leftarrow b$. Indeed, each of the two edges going into $a$ is coming from either a triangle or a rhombus; if both come from a triangle then the two triangles are touching. Thus, we can construct an infinite sequence of edges $a_{1}\leftarrow a_{2}\leftarrow a_{3}\leftarrow\dots$, so there must be a cycle in the graph, and we can perform a cycle flip, giving us a second tiling. ∎
@@ 5. GD Flips
Using the the idea from the proof of Lemma 4.1, we see that any two fine mixed subdivisions with the same arrangement of triangular holes differ by several cycle flips. Additionally, if we place $3$ unit triangles in the corners and one triangle of side n3n-3 in the middle we get an arrangement of triangular holes that admits only two fine mixed subdivisions, where only one cycle flip is possible. (See Figure 7.)
@@
For $n\geq 3$, consider the region obtained by removing three corner unit triangles from a triangle of side $n$, as well as removing a triangle of side n3n-3 in the middle. There are two configurations of rhombi that tile this region. Let the Clockwise GD (CW-GD) of size n3n-3 be the configuration where vertical rhombi (the ones without horizontal edges) go along the left side of the triangle, and let Counter-clockwise GD (CCW-GD) be the other one. [^1] This name is inspired by the (pre-2020) Google Drive logo. A GD flip consists of replacing a CW-GD with a CCW-GD. In particular, a GD flip is a cycle flip.
@@
First, we claim that there are no edges going from a rhombus strictly inside $C$ to a rhombus of $C$. Suppose there is such an edge $a\rightarrow b$. Without loss of generality, we may assume that the boundary between $a$ and $b$ is horizontal, that is, $a$ is directly below $b$. Then we can construct a sequence of rhombi starting with $b,$ $a$ where each next rhombus is directly below the previous one. Eventually, this sequence must hit $C$ again. Thus, we’ve constructed a path surrounded by $C$ connecting two vertices of $C$. This path creates a clockwise cycle smaller than $C$, which is a contradiction.
@@
Suppose that we begin with some tiling and start performing GD flips while possible. Since the number of vertical rhombi does not change after a GD flip, and the sum of $x$-coordinates of their centers strictly increases with every GD flip, we must eventually stop after finitely many flips. Note that there is at most one tiling with no clockwise cycles: if there were two, they would differ by several cycle flips, so one of them would have a clockwise cycle. This shows that starting from any tiling, we can get to the unique tiling with no clockwise cycles by performing GD flips, and then back to any other tiling by performing their inverses. ∎
@@ 7. Trapezoid Flips
For any triangle hole $s$, let $h_{1}(s),$ $h_{2}(s)$, and $h_{3}(s)$ denote the distances from sides $s$ to the parallel sides of the big triangle. Note that for any triangle $s$ with $h_{i}(s)>0$, there is a trapezoid flip that decreases $h_{i}(s)$ by $1$. Then the distance from a fine mixed subdivision with triangle set $S$ to the unique fine mixed subdivision with all triangles at the bottom is at most $\sum_{s\in S}h_{i}(s).$ Then the distance between two fine mixed subdivisions with triangle sets $S_{1}$ and $S_{2}$ is at most $\sum_{s\in S_{1}}h_{i}(s)+\sum_{s\in S_{2}}h_{i}(s).$ Since for any $s$ we have $h_{1}(s)+h_{2}(s)+h_{3}(s)=n-1,$ the sum of this expression over $i=1,2,3$ is equal to 2n(n1)2n(n-1), so for some choice of $i$ its at most $\frac{2n(n-1)}{3}$, as desired.
@@ 8. Further Questions
The rhombus tilings of a hexagon with sides $a,b,c,a,b,c$ are in bijection with $3$-dimensional Young tableaux that fit in a $a\times b\times c$ box. A GD flip of size 0 in this case is also known as a hexagon flip, and corresponds to removing a cube.
Figure 13. A GD flip of size 0 (note that this is rotated by $60^{\circ}$ counterclockwise relative to conventions in the rest of the paper for ease of visualization).